OpenStudy (kittykr98):

show how to solve the quadratic by factoring x^2 - 4x - 32 = 0 @studygurl14 @shelbygt520

OpenStudy (studygurl14):

to solve, find two numbers that when added together = -4 and when multiplied together, equals -32 First step, factor -32

OpenStudy (studygurl14):

Find teh factors of -32. Meaning the numbers that evenly go into -32 EXAMPLE: Factors of 6 are 1, 2, 3, and 6

OpenStudy (kittykr98):

so -2 + -2 = -4, is that a start?

OpenStudy (studygurl14):

-2*-2 don't equal -32

OpenStudy (studygurl14):

We're focusing on the factors of -32, remember? Not -4

OpenStudy (kittykr98):

oh ok my bad i was going by the original statement, "find two numbers that when added together equal -4"

OpenStudy (studygurl14):

That's the next step. :)

OpenStudy (kittykr98):

Ok so first find factors that when multiplied equal -32, thats the first step, right?

OpenStudy (kittykr98):

do the factors have to be the same two factors ?

OpenStudy (studygurl14):

yep. Sorry, I have to go eat dinner. b back tomorrow

OpenStudy (studygurl14):

and, no

OpenStudy (studygurl14):

I'll start you off. 1 and -31 are two factors. So are -1 and 32

OpenStudy (kittykr98):

so the factors would be -8 and 4 because -8*4=-32 and -8+4=-4

OpenStudy (kittykr98):

right? so whats the next step?

OpenStudy (anonymous):

can someone help me?

OpenStudy (kittykr98):

you have the same question?

jimthompson5910 (jim_thompson5910):

`so the factors would be -8 and 4 because -8*4=-32 and -8+4=-4` correct since the two numbers are -8 and 4, this means `x^2 - 4x - 32` factors to `(x-8)(x+4)`

OpenStudy (kittykr98):

and thats it? theres no other steps?

jimthompson5910 (jim_thompson5910):

so `x^2 - 4x - 32 = 0` becomes `(x-8)(x+4) = 0`

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):